OpenStudy (anonymous):

Given (x/a)+(y/b)=1 , the equation of the line t , determine a and b under the following conditions i) line t is 4 units from O, and the x-intercept is -8 ii) line t has equal intercepts and passes through (5, -3)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

Sorry, I can't think of the proper way to do these problems .-. I'll try tag some people for you tho

OpenStudy (sleepyhead314):

equation of a line HALP @Directrix @hartnn @freckles @satellite73 @Vocaloid

Directrix (directrix):

I have an idea for the second part. ii) line t has equal intercepts and passes through (5, -3) (x/a)+(y/b)=1 If the intercepts are equal, then a = b. The equation becomes: x/a + y/a = 1 The line passes through (5,-3) 5/a + (-3)/a = 1 Multiply through by a to get: 5 - 3 = a 2 = a The equation is: x/2 + y/2 = 1 which can be written as: y = 2 - x @AliLnn See what you think about this.

OpenStudy (anonymous):

o sorry i fall asleep... yea that looks like it

OpenStudy (sleepyhead314):

Thanks @Directrix :)